### v1

# Arens extensions of disjointness preserving multilinear operators on Riesz spaces

Abstract
Let $E_{1},\ldots,E_{m}$ be (non necessarily Archimedean) Riesz spaces, let $F$ be an Archimedean Riesz space and let $A\colon E_{1}\times\cdots\times E_{m}\to F$ be an order bounded disjointness preserving $m$-linear operator. We prove that all Arens extensions of $A$ are disjointness preserving if either $A$ has finite rank or the spaces are Banach lattices and $F^{*}$ has a Schauder basis consisting of disjointness preserving functionals.
1 Introduction
The study of disjointness preserving linear operators between Riesz spaces (vector lattices) started with the seminal work of Abramovich, Veskler and Koldunov [4] in 1979. Next, several authors investigated different aspects of such class of operators, see, e.g., [2, 3, 9, 12, 13, 19]. Multilinear disjointness preserving operators were considered later by Kusraev and Tabuev [15] in 2004, and then studied by Kusraev and Kusraeva in [14], by Toumi in [21], and, more recently, by Kusraeva in [16].
The main purpose of this work is to establish conditions so that Arens extensions of disjointness preserving multilinear operators between Riesz spaces are also disjointness preserving. The corresponding problems for Riesz multimorphisms, for almost Dunford-Pettis multilinear operators and for order continuous multilinear operators were addressed by the first two authors, respectively, in [6, 7, 8]. In our way to solve the aforementioned problem, we need first to study the linear case and to establish basic facts about disjointness preserving multilinear operators. This is done in Section 2, where, in particular, we: (i) Give conditions so that the second adjoint of a disjointess preserving linear operator $T\colon E\to F$ between Riesz spaces is also disjointness preserving. (ii) Prove that if $E_{1},\dots,E_{m}$ are Riesz spaces and $A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$ is a positive disjointess $m$-linear form, then there exist disjointness preserving positive linear functionals $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{m}\in E_{m}^{\sim}$ such that $A(x_{1},\ldots,x_{m})=\varphi_{1}(x_{1})\cdots\varphi_{m}(x_{m})$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. The main results, on disjointness preservation of Arens extensions of disjointness preserving multilinear operators, are proved in Section 3. It is important to mention that, in most of our results, the Riesz spaces on the domain of linear/multilinear operators are not supposed to be Archimedean as they were in [9, 14, 16, 19]. The case of operators between Banach lattices is also considered.
We refer the reader to [5, 18] for background on Riesz spaces and Banach lattices. Now, we present some terminology considering multilinear maps between Riesz spaces. Given Riesz spaces $E_{1},\ldots,E_{m},F$, the space of all order bounded (respectively, regular) $m$-linear operators from $E_{1}\times\cdots\times E_{m}$ to $F$ is denoted by ${\cal L}_{b}(E_{1},\ldots,E_{m};F)$ (respectively, ${\cal L}_{r}(E_{1},\ldots,E_{m};F)$). For precise definitions, see [14, 17]. When $F=\mathbb{R}$, we write ${\cal L}_{b}(E_{1},\ldots,E_{m})$ (respectively, ${\cal L}_{r}(E_{1},\ldots,E_{m})$). The order relation in these spaces is given by
$A\geq 0\Longleftrightarrow A(x_{1},\ldots,x_{m})\geq 0\mbox{~{}for all~{}}0\leq x_{1}\in E_{1},\ldots,0\leq x_{m}\in E_{m}.$
As usual, in the linear case $m=1$, the order dual of a Riesz space $E$ is denoted by $E^{\sim}$. The set of all permutations of $\{1,\ldots,m\}$ is denoted by $S_{m}$.
2 Disjointness preserving linear and multilinear operators
Recall that the (order) adjoint of an order bounded linear operator $T\colon E\to F$ between two Riesz spaces is the restriction of the adjoint of $T$ to $F^{\sim}$, denoted by $T^{\prime}\colon F^{\sim}\to E^{\sim}$. Of course, $T{{}^{\prime\prime}}:=(T^{\prime})^{\prime}$. As announced in the Introduction, we begin this section of preparatory results studying when the second adjoint $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ of a disjointness preserving linear operator $T\colon E\to F$ between two Riesz spaces is disjointness preserving as well. Let us recall that a linear operator $T\colon E\to F$ between two Riesz spaces is said to be disjointness preserving if $|T(x)|\wedge|T(y)|=0$ for all $x,y\in E$ with $|x|\wedge|y|=0$.
We begin our investigation considering order bounded linear operators between Riesz spaces. For a Riesz space $E$, by $J_{E}\colon E\to E^{\sim\sim}$ we denote the canonical Riesz homomorphism given by $J_{E}(x)(x^{\prime})=x^{\prime}(x)$ for $x\in E$ and $x^{\prime}\in E^{\sim}$.
Theorem 2.1.
Let $T\colon E\to F$ be an order bounded linear operator between two Riesz spaces.
(1) If $F$ is Archimedean and $T$ is disjointness preserving, then $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is disjointness preserving.
(2) If $J_{E}(E)=E^{\sim\sim}$ and $T$ is disjointness preserving, then $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is disjointness preserving.
(3) If $F^{\sim}$ separates the points of $F$ and $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is disjointness preserving, then $T$ is disjointness preserving.
Proof.
(1) By [5, Theorem 2.40] the modulus of $T$ exists and $|T|(|x|)=|T(|x|)|=|T(x)|$. Applying [5, Theorem 1.7] we get
$||T|(x)|$ $=||T|(x^{+})-|T|(x^{-})|=||T(x^{+})|-|T(x^{-})||$
$=|T(x^{+})+T(x^{-})|\wedge|T(x^{+})-T(x^{-})|=|T(|x|)|\wedge|T(x)|=|T|(|x|),$
hence $||T|(x)|=|T|(|x|)$ for every $x\in E$. This yields that $|T|$ is Riesz homomorphism, therefore the operator $|T|^{\prime\prime}$ is also a Riesz homomorphism by [5, Theorems 2.19 and 2.20]. Given $x^{\prime\prime},y^{\prime\prime}\in E^{\sim\sim}$ such that $x^{\prime\prime}\perp y^{\prime\prime}$, we have
00 0 $\leq|T^{\prime\prime}(x^{\prime\prime})|\wedge|T^{\prime\prime}(y^{\prime\prime})|\leq|T^{\prime\prime}|(|x^{\prime\prime}|)\wedge|T^{\prime\prime}|(|y^{\prime\prime}|)$
$\leq|T|^{\prime\prime}(|x^{\prime\prime}|)\wedge|T|^{\prime\prime}(|y^{\prime\prime}|)=|T|^{\prime\prime}(|x^{\prime\prime}|\wedge|y^{\prime\prime}|)=|T|^{\prime\prime}(0)=0,$
which implies that $T^{\prime\prime}(x^{\prime\prime})\perp T^{\prime\prime}(y^{\prime\prime})$, proving that $T^{\prime\prime}$ is disjointness prerserving.
(2) Since $T^{\prime\prime}(J_{E}(x))=J_{F}(T(x))$ for every $x\in E$, we get $T^{\prime\prime}(x)=J_{F}(T(x))$ for every $x\in J_{E}(E)=E^{\sim\sim}$, which gives that $T^{\prime\prime}$ is disjointness preserving because both $T$ and $J_{F}$ are disjointness preserving.
(3) Let $x,y\in E$ with $|x|\wedge|y|=0$. As $J_{E}\colon E\to E^{\sim\sim}$ is a Riesz hommorphism, we have $|J_{E}(x)|\wedge|J_{F}(y)|=0$, from which it follows that $|T^{\prime\prime}(J_{E}(x))|\wedge|T^{\prime\prime}(J_{E}(y))|=0$. Since $T^{\prime\prime}(J_{E}(z))=J_{F}(T(z))$ holds for every $z\in E$, we obtain that $|J_{F}(T(x))|\wedge|J_{F}(T(y))|=0$ in $F^{\sim\sim}$. Finally, since $F$ is a Riesz subspace of $F^{\sim\sim}$, $|T(x)|\wedge|T(y)|=0$ and we are done. ∎
If $F$ is a Banach lattice, then $F$ is Archimedean and its topological dual $F^{*}=F^{\sim}$ separates the points of $F$. Therefore, the following is a straightforward consequence of Theorem 2.1.
Corollary 2.2.
Let $E$ be a Riesz space and let $F$ be a Banach lattice. Then, an order bounded linear operator $T\colon E\to F$ is disjointness preserving if and only if its second adjoint $T^{\prime\prime}\colon E^{\sim\sim}\to F^{**}$ is disjointness preserving.
It is well known that a positive linear operator between Archimedean Riesz spaces is disjointness preserving if and only if it is a Riesz homomorphism (see, e.g., [19, p. 543]). Next we show that no condition on the underlying Riesz spaces is required for this to be true.
Proposition 2.3.
A positive linear operator $T\colon E\to F$ between Riesz spaces is disjointness preserving if and only if $T$ is a Riesz homomorphism. In this case, $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is also a Riesz homomorphism, hence disjointness preserving.
Proof.
Assuming that $T$ preserves disjointness, $|T(x^{+})|\wedge|T(x^{-})|=0$ holds for every $x\in E$. So, by [5, Theorem 1.7(5)],
$0=|T(x^{+})|\wedge|T(x^{-})|=\dfrac{1}{2}\big{|}|T(x^{+})+T(x^{-})|-|T(x^{+})-T(x^{-})|\big{|}=\dfrac{1}{2}\big{|}T(|x|)-|T(x)|\big{|}.$
Hence, $T(|x|)=|T(x)|$ for every $x\in E$, from which it follows that $T$ is a Riesz homomorphism. The converse is obvious. Calling on [5, Theorem 2.19 and 2.20], its second adjoint $T^{\prime\prime}$ is a Riesz homomorphism as well. Note that [5, Theorem 2.19] and the implication of [5, Theorem 2.20] we are using do not require any assumption on the spaces. In particular, $T^{\prime\prime}$ preserves disjointness. ∎
The following result, which is a multilinear version of [5, Theorem 1.14], shall be helpful in our study about disjointness preserving multilinear operators.
Theorem 2.4.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces with $F$ Archimedean and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be an $m$-linear operator. Suppose that, regardless of the positive vectors $x_{1}\in E_{1},\dots,x_{m}\in E_{m}$,
[…]
exists in $F$. Then the modulus of $A$ exists and $|A|(x_{1},\dots x_{m})$ coincides with the supremum above for all $x_{1}\in E_{1}^{+},\dots,x_{m}\in E_{m}^{+}.$
Proof.
We prove the bilinear case $m=2$. The general case follows analogously. By assumption, the map $B\colon E_{1}^{+}\times E_{2}^{+}\to F^{+}$ given by
$B(x_{1},x_{2})=\sup\left\{{|A(z_{1},z_{2})|}\,:\,{|z_{1}|\leq x_{1},\,|z_{2}|\leq x_{2}}\right\}$
is well defined. Reasoning as in the proof of [5, Theorem 1.14], $|A(z_{1},z_{2})|$ can be replace with $A(z_{1},z_{2})$ in the definition of $B$. Let us check now that $B$ is additive in each variable. Fix $y\in E_{2}^{+}$ and take $u,v\in E_{1}^{+}$. If $|u_{0}|\leq u$ and $|v_{0}|\leq v$, then $|u_{0}+v_{0}|\leq|u_{0}|+|v_{0}|\leq u+v$, which implies that
$A(u_{0},y)+A(v_{0},y)=A(u_{0}+v_{0},y)\leq B(u+v,y).$
A straightforward argument shows that $B(u,y)+B(v,y)\leq B(u+v,y)$. On the other hand, if $|w|\leq u+v$, there exits $u_{0},v_{0}\in E$ such that $|u_{0}|\leq u$, $|v_{0}|\leq v$ and $w=u_{0}+v_{0}$ (see [5, Theorem 1.13]). Hence,
$A(w,y)=A(u_{0},y)+A(v_{0},y)\leq B(u,y)+B(u,y),$
which gives that $B(u+v,w)\leq B(u,w)+B(v,w)$. Therefore $B(u+v,w)=B(u,w)+B(v,w)$. The same argument shows that $B$ is additive in the second variable as well. By [17, Theorem 2.3], $B$ can be extended to a positive bilinear map from $E_{1}\times E_{2}$ to $F$, which is also denote by $B$.
Now we prove that $B=\sup\{-A,A\}$. First, we observe that $-A\leq B$ and $A\leq B$ holds immediately. Let $T\colon E_{1}\times E_{2}\to F$ be a bilinear map such that $-A\leq T$ and $A\leq T$. It is clear that $T\geq 0$. Fix $x\in E_{1}^{+}$ and $y\in E_{2}^{+}$. Given $|u|\leq x$ and $|v|\leq y$, we have
$A(u,v)$ $=A(u^{+},v^{+})-A(u^{+},v^{-})-A(u^{-},v^{+})+A(u^{-},v^{-})$
$\leq T(u^{+},v^{+})+T(u^{+},v^{-})+T(u^{-},v^{+})+T(u^{-},v^{-})$
$=T(|u|,|v|)\leq T(x,y).$
Taking the supremum over all $|u|\leq x$ and $|v|\leq y$, we get $B(x,y)\leq T(x,y)$, which proves that $B=\sup\{-A,A\}=|A|$. ∎
From now on we focus on disjointness preserving multilinear operators. Given Riesz spaces $E_{1},\dots,E_{m},$ and $F$, an $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\to F$ is said to preserve disjointness if, for all $j\in\{1,\ldots,m\}$ and $x_{i}\in E_{i}^{+},\,i\neq j$, the linear operator
$A_{j}:=A(x_{1},\ldots,x_{j-1},\cdot,x_{j+1},\ldots,x_{m})\colon E_{j}\longrightarrow F,\,A_{j}(x_{j})=A(x_{1}\ldots,x_{m}),$
preserves disjointness.
Our first result on disjointness preserving multilinear operators is a multilinear counterpart of [5, Theorem 2.40].
Proposition 2.5.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces with $F$ Archimedean and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be an order bounded $m$-linear operator. If $A$ is disjointness preserving, then its modulus exists and
$|A(|x_{1}|,\ldots,|x_{m}|)|=||A|(x_{1},\ldots,x_{m})|=|A(x_{1}\ldots,x_{m})|=|A|(|x_{1}|,\ldots,|x_{m}|)$
for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. In particular, $|A|$ is disjointness preserving.
Proof.
Again, for simplicity, we prove the bilinear case $m=2$. We begin by noticing that, for all $a_{1}\in E_{1}^{+}$ and $a_{2}\in E_{2}^{+}$, the order bounded linear operators $A_{a_{1}}\colon E_{2}\longrightarrow F,A_{a_{1}}(x_{2})=A(a_{1},x_{2})$, and $A^{a_{2}}\colon E_{1}\longrightarrow F,A^{a_{2}}(x_{1})=A(x_{1},a_{2})$, preserve disjointness. So, by [5, Theorem 2.40] their modulus exist and
$|A_{a_{1}}|(|x_{2}|)=|A(a_{1},|x_{2}|)|=|A(a_{1},x_{2})|\text{ and }|A^{a_{2}}|(|x_{1}|)=|A(|x_{1}|,a_{2})|=|A(x_{1},a_{2})|$ (1)
for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Given $x_{1}\in E_{1}^{+}$, $x_{2}\in E_{2}^{+}$, $|z_{1}|\leq x_{1}$ and $|z_{2}|\leq x_{2}$, from (1) it follows that
$|A(z_{1},z_{2})|$ $=|A(z_{1}^{+},z_{2})-A(z_{1}^{-},z_{2})|\leq|A(z_{1}^{+},z_{2})|+|A(z_{1}^{-},z_{2})|$
$=|A_{z_{1}^{+}}|(|z_{2}|)+|A_{z_{1}^{-}}|(|z_{2}|)\leq|A_{z_{1}^{+}}|(x_{2})+|A_{z_{1}^{-}}|(x_{2})$
$=|A(z_{1}^{+},x_{2})|+|A(z_{1}^{-},x_{2})|=|A(z_{1}^{+},x_{2})+A(z_{1}^{-},x_{2})|$
$=|A(|z_{1}|,x_{2})|=|A^{x_{2}}|(|z_{1}|)$
$\leq|A^{x_{2}}|(x_{1})=|A(x_{1},x_{2})|,$
which implies that
$|A(x_{1},x_{2})|\leq\sup\left\{{|A(z_{1},z_{2})|}\,:\,{|z_{i}|\leq x_{i},\,i=1,2}\right\}\leq|A(x_{1},x_{2})|,$
proving that $\sup\left\{{|A(z_{1},z_{2})|}\,:\,{|z_{i}|\leq x_{i},\,i=1,2}\right\}=|A(x_{1},x_{2})|$ in $F$ for all $x_{1}\in E_{1}^{+}$ and $x_{2}\in E_{2}^{+}$. By Theorem 2.4, the modulus of $A$ exists and
$|A|(x_{1},x_{2})=\sup\left\{{|A(z_{1},z_{2})|}\,:\,{|z_{i}|\leq x_{i},\,i=1,2}\right\}=|A(x_{1},x_{2})|$
for all $x_{1}\in E_{1}^{+}$ and $x_{2}\in E_{2}^{+}$. In particular, $|A|(|x_{1}|,|x_{2}|)=|A(|x_{1}|,|x_{2}|)|$ holds for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Therefore, using $|A(x_{1},x_{2}^{+})|\wedge|A(x_{1},x_{2}^{-})|=0$ and (1),
$|A(x_{1},x_{2})|$ $=|A(x_{1},x_{2}^{+})-A(x_{1},x_{2}^{-})|=|A(x_{1},x_{2}^{+})-A(x_{1},x_{2}^{-})|=|A(x_{1},|x_{2}|)|=|A(|x_{1}|,|x_{2}|)|$
holds for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Then, for $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$,
$|A|(x_{1},x_{2})$ $=|A|(x_{1}^{+},x_{2}^{+})-|A|(x_{1}^{+},x_{2}^{-})-|A|(x_{1}^{-},x_{2}^{+})+|A|(x_{1}^{-},x_{2}^{-})$
$=|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{+})|+|A(x_{1}^{-},x_{2}^{-})|.$
Since $A$ is disjointness preserving, $|A(x_{1}^{+},x_{2}^{+})|\wedge|A(x_{1}^{-},x_{2}^{+})|=0$, so, by [5, Theorem 1.7(3)], we get
$\big{|}|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{-},x_{2}^{+})|\big{|}=|A(x_{1}^{+},x_{2}^{+})|\vee|A(x_{1}^{-},x_{2}^{+})|\leq|A|(|x_{1}|,x_{2}^{+}).$
Analogously, $\big{|}|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{-})|\big{|}\leq|A|(|x_{1}|,x_{2}^{-})$. Calling $a=|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{-},x_{2}^{+})|$ and $b=|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{-})|$, we obtain
$|a|\wedge|b|\leq|A|(|x_{1}|,x_{2}^{+})\wedge|A|(|x_{1}|,x_{2}^{-})=|A(x_{1},x_{2}^{+})|\wedge|A(x_{1},x_{2}^{-})|=0.$
By [5, Theorem 1.7],
$||A|(x_{1},x_{2})|$ $=|a-b|=|a+b|=|a|+|b|$
$=\big{|}|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{-},x_{2}^{+})|\big{|}+\big{|}|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{-})|\big{|}$
$=|A(x_{1}^{+},x_{2}^{+})|+|A(x_{1}^{-},x_{2}^{+})|+|A(x_{1}^{+},x_{2}^{-})|+|A(x_{1}^{-},x_{2}^{-})|$
$=|A|(x_{1}^{+},x_{2}^{+})+|A|(x_{1}^{-},x_{2}^{+})+|A|(x_{1}^{+},x_{2}^{-})+|A|(x_{1}^{-},x_{2}^{-})$
$=|A|(|x_{1}|,|x_{2}|),$
completing the proof of the desired formulas. The last statement follows from the formulas we have just proved. ∎
We conclude this preparatory section by proving that every disjointness preserving positive multilinear form can be written as the product of disjointness preserving positive linear functionals.
Theorem 2.6.
Let $E_{1},\ldots,E_{m}$ be Riesz spaces and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow\mathbb{R}$ be a positive $m$-linear form. If $A$ is disjointness preserving, then there exist disjointness preserving positive linear functionals $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{m}\in E_{m}^{\sim}$ such that $A(x_{1},\ldots,x_{m})=\varphi_{1}(x_{1})\cdots\varphi_{m}(x_{m})$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$.
Proof.
We proceed by induction on $m$. The result is immediate for $m=1$. Now we assume that the result holds for $m=k$ and we check its validity for $m=k+1$. To do so, let a positive disjointness preserving $(k+1)$-linear form $A\colon E_{1}\times\cdots\times E_{k+1}\longrightarrow\mathbb{R}$ be given. Assuming, without loss of generality, that $A\neq 0$, let $x_{0}^{1}\in E_{1}^{+},\ldots,x_{0}^{k}\in E_{k}^{+}$ be such that the map
$A_{x_{0}^{1},\ldots,x_{0}^{k}}\colon E_{k+1}\longrightarrow\mathbb{R}~{},~{}A_{x_{0}^{1},\ldots,x_{0}^{k}}(x_{k+1})=A(x_{0}^{1},\ldots,x_{0}^{k},x_{k+1}),$
is a non-zero disjointness preserving positive linear operator. Note that $A_{x_{0}^{1},\ldots,x_{0}^{k}}$ is a Riesz homomorphism by the proof of Theorem 2.1(1). And it is easy to check that
$A_{x_{1}+x_{0}^{1},\ldots,x_{k}+x_{0}^{k}}\geq A_{x_{0}^{1},\ldots,x_{0}^{k}}\geq 0\quad\text{and}\quad A_{x_{1}+x_{0}^{1},\ldots,x_{k}+x_{0}^{k}}\geq A_{x_{1},\ldots,x_{k}}\geq 0$
for all $x_{1}\in E_{1}^{+},\ldots,x_{k}\in E_{k}^{+}$. Applying Kutateladze’s Theorem [5, Theorem 2.50] for the inequalities above, there exist positive real numbers $t$ and $s$ such that
$A_{x_{0}^{1},\ldots,x_{0}^{k}}=tA_{x_{1}+x_{0}^{1},\ldots,x_{k}+x_{0}^{k}}\quad\text{and}\quad A_{x_{1},\ldots,x_{k}}=sA_{x_{1}+x_{0}^{1},\ldots,x_{k}+x_{0}^{k}}.$
As $A_{x_{0}^{1},\ldots,x_{0}^{k}}\neq 0$, $t\neq 0$, hence $A_{x_{1},\ldots,x_{k}}=\dfrac{s}{t}\,A_{x_{0}^{1},\ldots,x_{0}^{k}}$. This shows that, for all $x_{1}\in E_{1}^{+},\ldots,x_{k}\in E_{k}^{+}$, there exists $\gamma(x_{1},\ldots,x_{k})\in\mathbb{R}^{+}$ such that
$A_{x_{1},\ldots,x_{k}}=\gamma(x_{1},\ldots,x_{k})A_{x_{0}^{1},\ldots,x_{0}^{k}}.$
Taking $x_{0}^{k+1}\in E^{+}$ such that $A_{x_{0}^{1},\ldots,x_{0}^{k}}(x_{0}^{k+1})=1$, we get
$\gamma(x_{1},\dots,x_{k})=A_{x_{1},\dots,x_{k}}(x_{0}^{k+1})=A(x_{1},\dots,x_{k},x_{0}^{k+1})\text{~{}for all~{}}x_{1}\in E_{1}^{+},\dots,x_{k}\in E_{k}^{+}.$
Thus, the map $\gamma\colon E_{1}^{+}\times\cdots\times E_{k}^{+}\to\mathbb{R}^{+}$ is additive in each variable, and by [17, Theorem 2.3] it extends uniquely to a positive $k$-linear mapping $\widetilde{\gamma}\colon E_{1}\times\cdots\times E_{k}\to\mathbb{R}$ given by
$\widetilde{\gamma}(x_{1},\ldots,x_{k})=\sum_{{n_{i}=0,1}\atop{1\leq i\leq k}}(-1)^{\sum n_{i}}\gamma(x_{1n_{1}},\ldots,x_{kn_{k}}),$
where $x_{j}=x_{j0}-x_{j1},x_{j0}=x_{j}^{+},x_{j1}=x_{j}^{-}$. Let us see that $\widetilde{\gamma}$ is disjointness preserving. To do so, fix $i\in\{1,\dots,k\}$ and take $y_{i},\,z_{i}\in E_{i}$ such that $|y_{i}|\wedge|z_{i}|=0$. Given $x_{j}\in E_{j}^{+}$ for each $j\in\{1,\dots,k\}\setminus\{i\}$, we have
$|\widetilde{\gamma}(x_{1},\ldots,y_{i},\ldots,$ $x_{k})|\wedge|\widetilde{\gamma}(x_{1},\ldots,z_{i},\ldots,x_{k})|\leq$
$\leq\widetilde{\gamma}(x_{1},\ldots,|y_{i}|,\ldots,x_{k})\wedge\widetilde{\gamma}(x_{1},\ldots,|z_{i}|,\ldots,x_{k})$
$=\gamma(x_{1},\ldots,|y_{i}|,\ldots,x_{k})\wedge\gamma(x_{1},\ldots,|z_{i}|,\ldots,x_{k})$
$=A(x_{1},\ldots,|y_{i}|,\ldots,x_{k},x_{0}^{k+1})\wedge A(x_{1},\ldots,|z_{i}|,\ldots,x_{k},x_{0}^{k+1})=0,$
because $A$ is disjointness preserving. This proves that $\widetilde{\gamma}\colon E_{1}\times\cdots\times E_{k}\to\mathbb{R}$ is a disjointness preserving positive $k$-linear form. By the induction hypothesis, there exist positive disjointness preserving linear functionals $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{k}\in E_{k}^{\sim}$ such that $\widetilde{\gamma}(x_{1},\dots,x_{k})=\varphi_{1}(x_{1})\cdots\varphi_{k}(x_{k})$ for all $x_{1}\in E_{1},\dots,x_{k}\in E_{k}$. Finally, defining $\varphi_{k+1}=A_{x_{0}^{1},\dots,x_{0}^{k}}$, we have, for all $x_{1}\in E_{1},\ldots,x_{k+1}\in E_{k+1}$,
$A(x_{1},\ldots,x_{k},x_{k+1})$ $=\sum_{{n_{i}=0,1}\atop{1\leq i\leq k}}(-1)^{\sum n_{i}}A(x_{1n_{1}},\ldots,x_{kn_{k}},x_{k+1})$
$=\sum_{{n_{i}=0,1}\atop{1\leq i\leq k}}(-1)^{\sum n_{i}}A_{x_{1n_{1}},\ldots,x_{kn_{k}}}(x_{k+1})$
$=\sum_{{n_{i}=0,1}\atop{1\leq i\leq k}}(-1)^{\sum n_{i}}\gamma(x_{1n_{1}},\ldots,x_{kn_{k}})A_{x_{0}^{1},\ldots,x_{0}^{k}}(x_{k+1})$
$=\widetilde{\gamma}(x_{1},\ldots,x_{k})\varphi_{k+1}(x_{k+1})=\varphi_{1}(x_{1})\cdots\varphi_{k}(x_{k})\varphi_{k+1}(x_{k+1}),$
and we are done. ∎
3 Arens extensions of disjointness preserving multilinear operators
We shall use the notation from [7] to introduce the Arens extensions of multilinear operators. Let $m\in\mathbb{N}$ and $\rho\in S_{m}$ be given. For every $i=1,\ldots,m-1$, and $x_{\rho(i)}^{\prime\prime}\in E_{\rho(i)}^{\sim\sim}$, we consider the operator
$\overline{x_{\rho(i)}^{\prime\prime}}\colon\mathcal{L}_{b}(E_{\rho(i)},\ldots,E_{\rho(m)})\longrightarrow\mathcal{L}_{b}(E_{\rho(i+1)},\ldots,E_{\rho(m)})~{},~{}\overline{x_{\rho(i)}^{\prime\prime}}(B)=x_{\rho(i)}^{\prime\prime}\circ B^{i},$
where $B^{i}\colon E_{\rho(i+1)}\times\cdots\times E_{\rho(m)}\longrightarrow E_{\rho(i)}^{\sim}$ is given by
$B^{i}(x_{i+1},\ldots,x_{m})(x_{i})=B(x_{i},x_{i+1},\ldots,x_{m}),\text{ for all }x_{j}\in E_{\rho(j)},j=i,\ldots,m.$
In the case $i=m$, we define $\overline{x_{\rho(m)}^{\prime\prime}}\colon E_{\rho(m)}^{\sim}\longrightarrow\mathbb{R}$ by $\overline{x_{\rho(m)}^{\prime\prime}}=x_{\rho(m)}^{\prime\prime}$. The following lemma will be needed later.
Lemma 3.1.
For every $\rho\in S_{m}$, the operators $\overline{x_{\rho(1)}^{\prime\prime}},\dots,\overline{x_{\rho(m)}^{\prime\prime}}$ defined above are linear and order bounded. Moreover, if $0\leq x_{\rho(i)}^{\prime\prime}\in E_{i}^{\sim\sim}$ for some $i\in\{1,\dots,m\}$, then $\overline{x_{\rho(i)}^{\prime\prime}}$ is positive.
Proof.
It is clear that $\overline{x_{\rho(1)}^{\prime\prime}},\dots,\overline{x_{\rho(m)}^{\prime\prime}}$ are linear and that $\overline{x_{\rho(m)}^{\prime\prime}}$ is order bounded. Combining [17, Lemma 2.11] and [5, p. 15], we get $\mathcal{L}_{b}(E_{\rho(k)},\ldots,E_{\rho(m)})=\mathcal{L}_{r}(E_{\rho(k)},\ldots,E_{\rho(m)})$ for every $k=1,\ldots,m-1$. It follows from the proof of [6, Proposition 2.1] that $\overline{x_{\rho(i)}^{\prime\prime}}(B)$ is an order bounded operator for every $B\in\mathcal{L}_{b}(E_{\rho(i)},\ldots,E_{\rho(m)})$ and every $i=1,\dots,m-1$. The last statement follows immediately from [6, Proposition 2.1]. ∎
Given an $m$-linear form $C\colon E_{1}\times\cdots\times E_{m}\longrightarrow\mathbb{R}$, to each permutation $\rho\in S_{m}$ we consider the $m$-linear form
$C_{\rho}\colon E_{\rho(1)}\times\cdots\times E_{\rho(m)}\longrightarrow\mathbb{R}~{},~{}C_{\rho}(x_{1},\ldots,x_{m})=C(x_{\rho^{-1}(1)},\ldots,x_{\rho^{-1}(m)}).$
The Arens extension of an order bounded $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ with respect to the permutation $\rho$ is defined as the $m$-linear operator $AR_{m}^{\rho}(A)\colon E_{1}^{\sim\sim}\times\cdots\times E_{m}^{\sim\sim}\longrightarrow F^{\sim\sim}$ given by
$AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime})(y^{\prime})=\big{(}\overline{x_{\rho(m)}^{\prime\prime}}\circ\cdots\circ\overline{x_{\rho(1)}^{\prime\prime}}\big{)}((y^{\prime}\circ A)_{\rho})$
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\ldots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$ and $y^{\prime}\in F^{\sim}$. Although the notation is a bit different, these are exactly the Arens extensions studied in [6] and in many papers on Arens (or Aron-Berner) extensions. The extension $A^{(m+1)*}$ from [10, 11] is recovered by considering the permutation
$\theta(m)=1,\,\theta(m-1)=2,\ldots,\theta(2)=m-1,\,\theta(1)=m,$
that is, $AR_{m}^{\theta}(A)=A^{*(m+1)}$.
Let us see that the extensions $AR_{m}^{\rho}(A)$ are always order bounded.
Lemma 3.2.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces. If $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ is an $m$-linear order bounded operator, then $AR_{m}^{\rho}(A)$ is order bounded for every $\rho\in S_{m}$.
Proof.
Given $x_{1}^{\prime\prime}\in(E_{1}^{\sim\sim})^{+},\dots,x_{m-1}^{\prime\prime}\in(E_{m-1}^{\sim\sim})^{+}$, we need to prove that the map $B\colon E_{m}^{\sim\sim}\to F^{\sim\sim}$, defined by $B(x_{m}^{\prime\prime})=AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime})$, is order bounded. Indeed, if $O\subset E_{m}^{\sim\sim}$ is an order bounded set, then there exists $x^{\prime\prime},y^{\prime\prime}\in E_{m}^{\sim\sim}$ such that $x^{\prime\prime}\leq x_{m}^{\prime\prime}\leq y^{\prime\prime}$ for every $x_{m}^{\prime\prime}\in O$. Thus $|x_{m}^{\prime\prime}|\leq y^{\prime\prime}+2|x^{\prime\prime}|$ for every $x_{m}^{\prime\prime}\in O$ and
$|B(x_{m}^{\prime\prime})|$ $=|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},x_{m}^{\prime\prime})|\leq|AR_{m}^{\rho}(A)|(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},|x_{m}^{\prime\prime}|)$
$\leq|AR_{m}^{\rho}(A)|(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},y^{\prime\prime}+2|x^{\prime\prime}|)$
$\leq|AR_{m}^{\rho}(A)|(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},y^{\prime\prime})+2|AR_{m}^{\rho}(A)|(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},|x^{\prime\prime}|),$
which implies that $B$ is an order bounded operator, proving that $AR_{m}^{\rho}(A)$ is order bounded in the last variable. The same argument can be used to prove its order boundedness in the other variables. ∎
Our following result is an immediate consequence of Proposition 2.5.
Lemma 3.3.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces with $F$ Archimedean and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be an $m$-linear order bounded operator. Then, $A$ is disjointness preserving if and only if the linear operator $A_{j}\colon E_{j}\longrightarrow F,\,A_{j}(x_{j})=A(x_{1}\ldots,x_{m})$, is disjointness preserving for all $j\in\{1,\ldots,m\}$ and $x_{i}\in E_{i},\,i\neq j$.
The first result on Arens extensions that preserve disjointness for bilinear operators is the following.
Proposition 3.4.
Let $E_{1},E_{2},F$ be Riesz spaces with $F$ Archimedean and let $A\colon E_{1}\times E_{2}\longrightarrow F$ be an order bounded bilinear operator. If $A$ is disjointness preserving, then for each $\rho\in S_{2}$ and $x_{1}\in E_{1},x_{2}\in E_{2}$, the operators
$AR_{2}^{\rho}(A)(J_{E_{1}}(x_{1}),\cdot)\colon E_{2}^{\sim\sim}\longrightarrow F^{\sim\sim},\,AR_{2}^{\rho}(A)(J_{E_{1}}(x_{1}),\cdot)(y_{2}^{\prime\prime})=AR_{2}^{\rho}(A)(J_{E_{1}}(x_{1}),y_{2}^{\prime\prime})$
and
$AR_{2}^{\rho}(A)(\cdot,J_{E_{2}}(x_{2}))\colon E_{1}^{\sim\sim}\longrightarrow F^{\sim\sim},\,AR_{2}^{\rho}(A)(\cdot,J_{E_{2}}(x_{2}))(x_{2}^{\prime\prime})=AR_{2}^{\rho}(A)(x_{1}^{\prime\prime},J_{E_{2}}(x_{2})),$
are disjointness preserving.
Proof.
In this case we only have two permutations, namely $\theta$ ($\theta(1)=2,\theta(2)=1$) and $id$ ($id(1)=1,id(2)=2)$. Let us check that $AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1}),\cdot)=A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving for every $x_{1}\in E_{1}$. Indeed, for all $x_{1}\in E_{1},\,x_{2}^{\prime\prime}\in E_{2}^{\sim\sim}$ and $y^{\prime}\in F^{\sim}$,
$A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),x_{2}^{\prime\prime})(y^{\prime})$ $=J_{E_{1}}(x_{1})\big{(}\overline{x_{2}^{\prime\prime}}^{\theta}(y^{\prime}\circ A)\big{)}=\overline{x_{2}^{\prime\prime}}^{\theta}(y^{\prime}\circ A)(x_{1})$
$=x_{2}^{\prime\prime}(y^{\prime}\circ A(x_{1},\cdot))=x_{2}^{\prime\prime}(A(x_{1},\cdot)^{\prime}(y^{\prime}))=A(x_{1},\cdot)^{\prime\prime}(x_{2}^{\prime\prime})(y^{\prime}),$
hence $A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),x_{2}^{\prime\prime})=A(x_{1},\cdot)^{\prime\prime}(x_{2}^{\prime\prime})$ for all $x_{1}\in E_{1}$ and $x_{2}^{\prime\prime}\in E_{2}^{\sim\sim}$. Since $A(x_{1},\cdot)$ is disjointness preserving by Lemma 3.3, so is its second adjoint $A(x_{1},\cdot)^{\prime\prime}$ by Theorem 2.1(a). Therefore, $A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving. The same argument shows that $AR_{2}^{id}(A)(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving.
Now, we check that $AR_{2}^{\theta}(A)(\cdot,J_{E_{2}}(x_{2}))=A^{\ast\ast\ast}(\cdot,J_{E_{2}}(x_{2}))$ is also disjointness preserving. First, observe that for all $x_{1}\in E_{1},x_{2}\in E_{2}$ and $y^{\prime}\in F^{\sim}$,
$\overline{J_{E_{2}}(x_{2})}^{\theta}(y^{\prime}\circ A)(x_{1})=J_{E_{2}}(x_{2})(y^{\prime}\circ A(x_{1},\cdot))=y^{\prime}(A(x_{1},x_{2}))=[y^{\prime}\circ A(\cdot,x_{2})](x_{1}),$
which implies that $\overline{J_{E_{2}}(x_{2})}^{\theta}(y^{\prime}\circ A)=y^{\prime}\circ A(\cdot,x_{2})$. Thus, given $x_{1}^{\ast\ast}\in E_{1}^{\ast\ast}$, $x_{2}\in E_{2}$ and $y^{\prime}\in F^{\sim}$, we have
$A^{\ast\ast\ast}(x_{1}^{\prime\prime},J_{E_{2}}(x_{2}))(y^{\prime})$ $=x_{1}^{\prime\prime}\big{(}\overline{J_{E_{2}}(x_{2})}^{\theta}(y^{\prime}\circ A)\big{)}=x_{1}^{\prime\prime}(y^{\prime}\circ A(\cdot,x_{2}))$
$=x_{1}^{\prime\prime}(A(\cdot,x_{2}))^{\prime}(y^{\prime}))=A(\cdot,x_{2})^{\prime\prime}(x_{1}^{\prime\prime})(y^{\prime}).$
This proves that $A^{\ast\ast\ast}(x_{1}^{\prime\prime},J_{E_{2}}(x_{2}))=A(\cdot,x_{2})^{\prime\prime}(x_{1}^{\prime\prime})$ for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim}$ and $x_{2}\in E_{2}$. Since $A(\cdot,x_{2})$ is disjointness preserving by Lemma 3.3, we get from Theorem 2.1(a) that its second adjoint is also disjointness preserving, which implies that $A^{\ast\ast\ast}(\cdot,J_{E_{2}}(x_{2}))$ is disjointness preserving as well. The same argument shows that $AR_{2}^{id}(A)(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving. ∎
As in [6, p. 11], a map taking values in a Riesz space is said to have finite rank if the sublattice generated by its range is finite dimensional. Now, we prove our first main result, which shows that all Arens extensions of a finite rank order bounded disjointness preserving multilinear operator into an Archimedean Riesz space are also disjointness preserving.
Theorem 3.5.
Let $E_{1},\ldots,E_{m},F$ be Riesz space with $F$ Archimedean and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a finite rank order bounded $m$-linear operator. If $A$ is disjointness preserving, then all Arens extensions $AR_{m}^{\rho}(A)$ of $A$, $\rho\in S_{m}$, are disjointness preserving.
Proof.
We split the proof into three steps. First we prove the case $F=\mathbb{R}$. Given an order bounded disjointness preserving $m$-linear form $A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$, we obtain from Proposition 2.5 and Theorem 2.6 that there are disjointness preserving linear functionals $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{m}\in E_{m}^{\sim}$ such that $|A|(x_{1},\ldots,x_{m})=\varphi_{1}(x_{1})\cdots\varphi_{m}(x_{m})$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. By the proof of [6, Theorem 3.2],
$AR_{m}^{\rho}(|A|)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime})=\varphi_{1}^{\prime\prime}(x_{1}^{\prime\prime})\cdots\varphi_{m}^{\prime\prime}(x_{m}^{\prime\prime})$
holds for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\ldots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$ and $\rho\in S_{m}$. Since $\varphi_{1},\dots,\varphi_{m}$ are disjointness preserving, their second adjoints $\varphi_{1}^{\prime\prime},\dots,\varphi_{1}^{\prime\prime}$ are disjointness preserving by Theorem 2.1(a). Then, $AR_{m}^{\rho}(|A|)$ is disjointness preserving. Using Lemma 3.1 and the fact that the composition of positive operators is a positive operator, it is easy to check that the correspondence
$A\in\mathcal{L}_{b}(E_{1},\ldots,E_{m};F)\mapsto AR_{m}^{\rho}(A)\in\mathcal{L}_{b}(E_{1}^{\sim\sim},\ldots,E_{m}^{\sim\sim};F^{\sim\sim})$
is a positive linear operator. So, $A\leq|A|$ implies $AR_{m}^{\rho}(A)\leq AR_{m}^{\rho}(|A|)$, from which it follows that $AR_{m}^{\rho}(A)$ is also disjointness preserving.
In the case $F=\mathbb{R}^{n},n\in\mathbb{N}$, we apply the previous case to the compositions $\pi_{j}\circ A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}^{n}$, where $\pi_{j}\colon\mathbb{R}^{n}\to\mathbb{R}$ is the projection in the $j$-th coordinate, $j=1,\ldots,n$, to obtain that $AR_{m}^{\rho}(\pi_{j}\circ A)\colon E_{1}^{\sim\sim}\times\cdots\times E_{m}^{\sim\sim}\longrightarrow\mathbb{R}$ is disjointness preserving for all $\rho\in S_{m}$ and $j=1,\dots,n$. By [6, Remark 2.3],
$\pi_{j}\circ AR_{m}^{\rho}(A)=\pi_{j}^{\prime\prime}\circ AR_{m}^{\rho}(A)=AR_{m}^{\rho}(\pi_{j}\circ A)$
holds for all $\rho\in S_{m}$ and $j=1,\dots,m$. Therefore,
$AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime})$ $=(\pi_{1}\circ AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime}),\dots,\pi_{n}\circ AR_{m}^{\rho}(A))(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime}))$
$=(AR_{m}^{\rho}(\pi_{1}\circ A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime}),\dots,AR_{m}^{\rho}(\pi_{n}\circ A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime}))$
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$. Considering $\mathbb{R}^{n}$ with the coordinatewise order, implies that $AR_{m}^{\rho}(A)$ is disjointness preserving for every $\rho\in S_{m}$.
Finally, let $F$ be an Archimedean Riesz space $F$ and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a finite rank order bounded disjointness preserving operator. Calling $G$ be the finite-dimensional sublattice of $F$ generated by the range of $A$, by $i\colon G\longrightarrow F$ we denote the inclusion operator. It is clear that $A_{1}\colon E_{1}\times\cdots\times E_{m}\longrightarrow G,A_{1}(x_{1},\ldots,x_{m})=A(x_{1},\ldots,x_{m})$, is a disjointness preserving $m$-linear operator. Since $G$ is Archimedean, by [20, Corollary 1, p. 70] there exists a Riesz isomorphism $I\colon G\to\mathbb{R}^{n}$, where $n=\dim G$. Thus, $I\circ A_{1}$ is an $\mathbb{R}^{n}$-valued disjointness preserving multilineara operator, so, by the second part of this proof, we obtain that $AR_{m}^{\rho}(I\circ A_{1})$ is disjointness preserving. From [6, Remark 2.3] we have
$AR_{m}^{\rho}(A)=AR_{m}^{\rho}(i\circ I^{-1}\circ I\circ A_{1})=i^{\prime\prime}\circ(I^{-1})^{\prime\prime}\circ AR_{m}^{\rho}(I\circ A_{1}).$
Therefore, $AR_{m}^{\rho}(A)$ is disjointness preserving because $i^{\prime\prime}$ and $(I^{-1})^{\prime\prime}$ are disjointness preserving linear operators by Theorem 2.4. ∎
Now we proceed to the case of arbitrary order bounded multilinear operators. The following two results shall be helpful.
Proposition 3.6.
Let $E_{1},\dots,E_{m},\,F$ be Riesz spaces, let $A\colon E_{1}\times\cdots\times E_{m}\to F$ be an order bounded disjointness preserving $m$-linear map, and let $\rho\in S_{m}$ be given. If $y^{\prime}\in F^{\sim}$ is disjointness preserving, then $J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ is disjointness preserving and
$|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|(|y^{\prime}|)=|AR_{m}^{\rho}(A)(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)(y^{\prime})|=|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})(y^{\prime})|$
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$.
Proof.
If $y^{\prime}\in F^{\sim}$ is disjointness preserving, then $y^{\prime}\circ A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$ is an order bounded $m$-linear disjointness preserving operator. By Theorem 3.5 we get that $AR_{m}^{\rho}(y^{\prime}\circ A)$ is also disjointness preserving. The proof of [6, Proposition 3.4] yields that $AR_{m}^{\rho}(y^{\prime}\circ A)=J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ holds, proving that $J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ is disjointness preserving. Calling on Proposition 2.5,
$|J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)|=|J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|$ (2)
holds for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$. Observe also that $J_{F^{\sim}}(y^{\prime})$ is disjointness preserving by Theorem 2.1(1), hence
$|J_{F^{\sim}}(y^{\prime})|(|y^{\prime\prime}|)=|J_{F^{\sim}}(y^{\prime})(y^{\prime\prime})|$ (3)
for all $y^{\prime\prime}\in F^{\sim\sim}$. Combining (2) and (3), we have
$|AR_{m}^{\rho}(A)(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)(y^{\prime})|$ $=|(J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A))(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)|$
$=|(J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A))(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|$
$=|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})(y^{\prime})|$
$\leq|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|(|y^{\prime}|)$
$=J_{F^{\sim}}(|y^{\prime}|)(|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|)$
$=|J_{F^{\sim}}(y^{\prime})|(|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|)$
$=|J_{F^{\sim}}(y^{\prime})(AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime}))|$
$=|AR_{m}^{\rho}(A)(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)(y^{\prime})|$
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$. ∎
Next result can be regarded as a version of [6, Proposition 3.6] for disjointness preserving multilinear operators.
Proposition 3.7.
Let $E_{1},\ldots,E_{m},F$ be Banach lattices, let $A\in{\cal L}_{b}(E_{1},\ldots,E_{m};F)$ be a disjointness preserving operator and let $\rho\in S_{m}$ be given. Then:
(a) $y^{***}\circ AB_{m}^{\rho}(A)$ is a disjointness preserving $m$-linear operator for every weak*-continuous disjointness preserving functional $y^{***}\in F^{***}$.
(b) Fix $i\in\{1,\ldots,m\}$ and let $w_{i}^{\ast\ast},z_{i}^{\ast\ast}\in E_{i}^{\ast\ast}$ be such that $|w_{i}^{\ast\ast}|\wedge|z_{i}^{\ast\ast}|=0$. Then
$(|AB_{m}^{\rho}(A)(x_{1}^{**},\ldots,w_{i}^{\ast\ast},\ldots,x_{m}^{**})|\wedge|AB_{m}^{\rho}(A)(x_{1}^{**},\ldots,z_{i}^{\ast\ast},\ldots,x_{m}^{**})|)(y^{*})=0$
for all $x_{j}^{**}\in E_{j}^{**},j\neq i$ and any $y^{*}\in\overline{{\rm span}}\{\varphi\in F^{\ast}:\varphi\text{ is disjointness preserving}\}.$
Proof.
(a) Let $y^{***}\in F^{***}$ be a weak*-continuous disjointness preserving functional. Since $y^{***}$ is weak*-continuous, there exists $y^{\ast}\in F^{\ast}$ such that $y^{\ast\ast\ast}=J_{F^{\ast}}(y^{\ast})$. As $y^{\ast\ast\ast}$ is disjointness preserving and and $J_{F}$ is a Riesz homomorphisms, $y^{\ast}=y^{\ast\ast\ast}\circ J_{F}$ is disjointness preserving. Now, Proposition 3.6 yields that $y^{***}\circ AB_{m}^{\rho}(A)=J_{F^{\ast}}(y^{\ast})\circ AB_{m}^{\rho}(A)$ is a disjointness preserving operator.
(b) For each $j\neq i$, take $x_{j}^{\ast\ast}\in E_{j}^{\ast\ast}$. Given $y^{\ast}\in\overline{{\rm span}}\{\varphi\in F^{\ast}:\varphi\text{ is a disjointness preserving}\}$, there exists a sequence $(y_{n}^{\ast})_{n=1}^{\infty}$ in ${\rm span}\{\varphi\in F^{\ast}:\varphi\text{ is a disjointness preserving}\}$ such that $y_{n}^{\ast}\longrightarrow y^{\ast}$, hence $|y_{n}^{\ast}|\longrightarrow|y^{\ast}|$. Moreover, for each $n\in\mathbb{N}$ there exist disjointness preserving linear functionals $\varphi^{1}_{n},\ldots,\varphi_{n}^{k_{n}}\in F^{\ast}$ and scalars $\alpha_{n}^{1},\ldots,\alpha_{n}^{k_{n}}$ such that $y_{n}^{\ast}=\sum\limits_{j=1}^{k_{n}}\alpha_{n}^{j}\varphi_{n}^{j}$. For each $w\in E_{i}^{\ast\ast}$, write
$AB_{m}^{\rho}(A)_{i}(w):=AB_{m}^{\rho}(A)(x_{1}^{\ast\ast},\ldots,x_{i-1}^{\ast\ast},w,x_{i+1}^{\ast\ast},\ldots,x_{m}^{\ast\ast}).$
It follows from Proposition 3.6 that $|AB_{m}^{\rho}(A)_{i}(w)|(|\varphi|)=|AB_{m}^{\rho}(A)_{i}(w)(\varphi)|$ for every disjointness preserving $\varphi\in F^{\ast}$ and every $w\in E_{i}^{\ast\ast}$. Therefore,
$\big{(}|AB_{m}^{\rho}$ $(A)_{i}(w_{i}^{\ast\ast})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|\big{)}(|y^{*}|)$
$=\lim_{n\rightarrow\infty}\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|\big{)}(|y_{n}^{\ast}|)$
$=\lim_{n\rightarrow\infty}\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|\big{)}\Big{(}\Big{|}\sum\limits_{j=1}^{k_{n}}\alpha_{n}^{j}\varphi_{n}^{j}\Big{|}\Big{)}$
$\leq\lim_{n\rightarrow\infty}\displaystyle\sum_{j=1}^{k_{n}}|\alpha_{n}^{j}|\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|\big{)}(|\varphi_{n}^{j}|)$
$\leq\lim_{n\rightarrow\infty}\displaystyle\sum_{j=1}^{k_{n}}|\alpha_{n}^{j}|\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})|(|\varphi_{n}^{j}|)\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|(|\varphi_{n}^{j}|)\big{)}$
$=\lim_{n\rightarrow\infty}\displaystyle\sum_{j=1}^{k_{n}}|\alpha_{n}^{j}|\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})(\varphi_{n}^{j})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})(\varphi_{n}^{j})|\big{)}$
$=\lim_{n\rightarrow\infty}\displaystyle\sum_{j=1}^{k_{n}}|\alpha_{n}^{j}|\big{(}|(J_{F^{\ast}}(\varphi_{n}^{j})\circ AB_{m}^{\rho}(A)_{i})(w_{i}^{\ast\ast})|\wedge|(J_{F^{\ast}}(\varphi_{n}^{j})\circ AB_{m}^{\rho}(A)_{i})(z_{i}^{\ast\ast})|\big{)}=0,$
where the last equality follows from Proposition 3.6. ∎
Our second main result is a consequence of Proposition 3.7(b).
Corollary 3.8.
Let $F$ be a Banach lattice such that $F^{\ast}$ has a Schauder basis consisting of disjointness preserving linear functionals. Then all Arens extensions of any $F$-valued disjointness preserving multilinear operator are disjointness preserving.
Example 3.9.
Corollary 3.8 holds for the following choices of $F$: $c_{0}$, $\ell_{p}$ with $1<p<\infty$, Tsirelson’s original space $T^{\ast}$ and its dual $T$, Schreier’s space $S$, and the predual $d_{\ast}(w,1)$ of the Lorentz sequence space $d(w,1)$ (see [6, Example 3.9]).