### v2

# Arens extensions of disjointness preserving multilinear operators on Riesz spaces

Abstract
Let $E_{1},\ldots,E_{m}$ be (non necessarily Archimedean) Riesz spaces, let $F$ be a Dede-kind complete Riesz space and let $A\colon E_{1}\times\cdots\times E_{m}\to F$ be a regular disjointness preserving $m$-linear operator. We prove that all Arens extensions of $A$ are disjointness preserving if either $A$ has finite rank or the spaces are Banach lattices and $F^{*}$ has a Schauder basis consisting of disjointness preserving functionals.
1 Introduction
The study of disjointness preserving linear operators between Riesz spaces (vector lattices) started with the seminal work of Abramovich, Veskler and Koldunov [4] in 1979. Next, several authors investigated different aspects of such class of operators, see, e.g., [2, 3, 9, 14, 15, 23]. Multilinear disjointness preserving operators were considered later by Kusraev and Tabuev [18] in 2004, and then studied by Bu, Buskes and Kusraev in [11], by Kusraev and Kusraeva in [16], by Toumi in [25], and, more recently, by Kusraeva in [19, 20].
The main purpose of this work is to establish conditions so that Arens extensions of disjointness preserving multilinear operators between Riesz spaces are also disjointness preserving. The corresponding problems for Riesz multimorphisms, for almost Dunford-Pettis multilinear operators and for order continuous multilinear operators were addressed by the first two authors, respectively, in [6, 7, 8]. In our way to solve the aforementioned problem, we need first to study the linear case and to establish basic facts about disjointness preserving multilinear operators. This is done in Section 2, where, in particular, we give conditions so that the second adjoint of a disjointess preserving linear operator $T\colon E\to F$ between Riesz spaces is also disjointness preserving. Our main results, announced in the Abstract, on Arens extensions of disjointness preserving multilinear operators between Riesz spaces/Banach lattices are proved in Section 3. It is important to mention that, in most of our results, the Riesz spaces on the domain of linear/multilinear operators are not supposed to be Archimedean as they were in [9, 16, 19, 23].
We refer the reader to [5, 22] for background on Riesz spaces and Banach lattices. Now, we present some terminology considering multilinear maps between Riesz spaces. Given Riesz spaces $E_{1},\ldots,E_{m},F$, the space of all regular $m$-linear operators from $E_{1}\times\cdots\times E_{m}$ to $F$ is denoted by ${\cal L}_{r}(E_{1},\ldots,E_{m};F)$. For precise definitions, see [21]. When $F=\mathbb{R}$, we write ${\cal L}_{r}(E_{1},\ldots,E_{m})$. The order relation in this space is given by
$A\geq 0\Longleftrightarrow A(x_{1},\ldots,x_{m})\geq 0\mbox{~{}for all~{}}0\leq x_{1}\in E_{1},\ldots,0\leq x_{m}\in E_{m}.$
The positive cone of a Riesz space $E$ is denoted by $E^{+}$. As usual, in the linear case $m=1$, the order dual of a Riesz space $E$ is denoted by $E^{\sim}$ and the topological dual of a Banach lattice $E$ by $E^{*}$. The symbol $x\perp y$ means that the vectors $x$ and $y$ of a Riesz space are disjoint. The set of all permutations of $\{1,\ldots,m\}$ is denoted by $S_{m}$.
2 Disjointness preserving linear and multilinear operators
Recall that the (order) adjoint of an order bounded linear operator $T\colon E\to F$ between two Riesz spaces is the restriction of the adjoint of $T$ to $F^{\sim}$, denoted by $T^{\prime}\colon F^{\sim}\to E^{\sim}$. Of course, $T{{}^{\prime\prime}}:=(T^{\prime})^{\prime}$. As announced in the Introduction, we begin this section of preparatory results studying when the second adjoint $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ of a disjointness preserving linear operator $T\colon E\to F$ between two Riesz spaces is disjointness preserving as well. Let us recall that a linear operator $T\colon E\to F$ between two Riesz spaces is said to be disjointness preserving if $T(x)\perp T(y)$ in $F$ whenever $x\perp y$ in $E$.
We begin our investigation considering order bounded linear operators between Riesz spaces. For a Riesz space $E$, by $J_{E}\colon E\to E^{\sim\sim}$ we denote the canonical Riesz homomorphism given by $J_{E}(x)(x^{\prime})=x^{\prime}(x)$ for $x\in E$ and $x^{\prime}\in E^{\sim}$.
Theorem 2.1.
Let $T\colon E\to F$ be an order bounded linear operator between two Riesz spaces.
(1) If $F$ is Archimedean and $T$ is disjointness preserving, then $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is disjointness preserving.
(2) If $J_{E}(E)=E^{\sim\sim}$ and $T$ is disjointness preserving, then $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is disjointness preserving.
(3) If $F^{\sim}$ separates the points of $F$ and $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is disjointness preserving, then $T$ is disjointness preserving.
Proof.
(1) By [5, Theorem 2.40] the modulus of $T$ exists and $|T|(|x|)=|T(|x|)|=|T(x)|$ for every $x\in E$. Applying [5, Theorem 1.7] we get
$||T|(x)|$ $=||T|(x^{+})-|T|(x^{-})|=||T(x^{+})|-|T(x^{-})||$
$=|T(x^{+})+T(x^{-})|\wedge|T(x^{+})-T(x^{-})|=|T(|x|)|\wedge|T(x)|=|T|(|x|),$
hence $||T|(x)|=|T|(|x|)$ for every $x\in E$. This yields that $|T|$ is Riesz homomorphism, therefore the operator $|T|^{\prime\prime}$ is also a Riesz homomorphism by [5, Theorems 2.19 and 2.20]. Given $x^{\prime\prime},y^{\prime\prime}\in E^{\sim\sim}$ such that $x^{\prime\prime}\perp y^{\prime\prime}$, we have
00 0 $\leq|T^{\prime\prime}(x^{\prime\prime})|\wedge|T^{\prime\prime}(y^{\prime\prime})|\leq|T^{\prime\prime}|(|x^{\prime\prime}|)\wedge|T^{\prime\prime}|(|y^{\prime\prime}|)$
$\leq|T|^{\prime\prime}(|x^{\prime\prime}|)\wedge|T|^{\prime\prime}(|y^{\prime\prime}|)=|T|^{\prime\prime}(|x^{\prime\prime}|\wedge|y^{\prime\prime}|)=|T|^{\prime\prime}(0)=0,$
which implies that $T^{\prime\prime}(x^{\prime\prime})\perp T^{\prime\prime}(y^{\prime\prime})$, proving that $T^{\prime\prime}$ is disjointness prerserving.
(2) Since $T^{\prime\prime}(J_{E}(x))=J_{F}(T(x))$ for every $x\in E$, we get $T^{\prime\prime}(x)=J_{F}(T(x))$ for every $x\in J_{E}(E)=E^{\sim\sim}$, which gives that $T^{\prime\prime}$ is disjointness preserving because both $T$ and $J_{F}$ are disjointness preserving.
(3) Let $x$ and $y$ be disjoint vectors in $E$. As $J_{E}\colon E\to E^{\sim\sim}$ is a Riesz homomorphism, we have $J_{E}(x)\perp J_{F}(y)$, from which it follows that
$J_{F}(T(x))=T^{\prime\prime}(J_{E}(x))\perp T^{\prime\prime}(J_{E}(y))=J_{F}(T(y))\mbox{ in }F^{\sim\sim}$
because $T^{\prime\prime}$ preserves disjointness. Using now that $(J_{F})^{-1}\colon J_{F}(F)\to F$ is a Riesz homomorphism we get $T(x)\perp T(y)$. ∎
If $F$ is a Banach lattice, then $F$ is Archimedean and its topological dual $F^{*}=F^{\sim}$ separates the points of $F$. Therefore, the following is a straightforward consequence of Theorem 2.1.
Corollary 2.2.
Let $E$ be a Riesz space and let $F$ be a Banach lattice. Then, an order bounded linear operator $T\colon E\to F$ is disjointness preserving if and only if its second adjoint $T^{\prime\prime}\colon E^{\sim\sim}\to F^{**}$ is disjointness preserving.
Remark 2.3.
The situation is quite simpler for positive operators. The following are equivalent for a positive linear operator $T\colon E\to F$ between Riesz spaces.
(a) $T$ is disjointness preserving.
(b) $T$ is a Riesz homomorphism.
(c) $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is disjointness preserving.
(d) $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is a Riesz homomorphism.
Indeed, (a)$\Leftrightarrow$(b) is well known (and it is easy to be checked directly); so (c)$\Leftrightarrow$(d) also holds because $T^{\prime\prime}\geq 0$. (a)$\Rightarrow$(d) follows from a combination of [5, Theorem 2.19], which imposes no conditions on $E$ and $F$, with the implication of [5, Theorem 2.20] that does not require any assumption on the spaces. For (d)$\Rightarrow$(a), use that $E$ and $F$ are (Riesz isomorphic to) sublattices of $E^{\sim\sim}$ and $F^{\sim\sim}$, or, alternatively, that $T=(J_{F})^{-1}\circ T^{\sim\sim}\circ J_{E}$, $(J_{F})^{-1}\colon J_{F}(F)\to F$.
From now on we focus on disjointness preserving multilinear operators, a topic for which we gave several references in the Introduction. Given Riesz spaces $E_{1},\dots,E_{m},$ and $F$, an $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\to F$ is said to preserve disjointness if $A$ preserves disjointness separately for fixed positive vectors, that is, for all $j\in\{1,\ldots,m\}$ and $x_{i}\in E_{i}^{+},\,i\neq j$, the linear operator
$x_{j}\in E_{j}\longrightarrow A(x_{1}\ldots,x_{m})\in F$
preserves disjointness. It is not difficult to check that this is equivalent to say that $A$ preserves disjointness separately, meaning that the above holds for all $j\in\{1,\ldots,m\}$ and $x_{i}\in E_{i},i\neq j$.
It is stated in [16, Theorem 3] that, for every order bounded disjointness preserving $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\to F$ between Archimedean Riesz spaces, $|A|$ exists and satisfies
$|A(|x_{1}|,\ldots,|x_{m}|)|=||A|(x_{1},\ldots,x_{m})|=|A(x_{1}\ldots,x_{m})|=|A|(|x_{1}|,\ldots,|x_{m}|)$
for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. Since [16, Theorem 3] is just a restatement of a result in the difficult-to-read reference [18] (it is in Russian), we prove next a similar version for regular operators that shall fit our purposes.
Recall that an $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\to F$ between Riesz spaces is a Riesz multimorphism (or an $m$-morphism) if $|A(x_{1},\ldots,x_{m})|=A(|x_{1}|,\ldots,|x_{m}|)$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. Or, equivalently, if $A$ is a Riesz homomorphism separately for fixed positive vectors. As in the linear case, a positive multilinear operator preserves disjointness if and only if it is a Riesz multimorphism.
A positive partition of a positive vector $x$ of a Riesz space $E$ is a finite sequence of elements of $E^{+}$ whose sum equals $x$. The set of all positive partitions of $x$ shall be denoted by $\prod_{x}$.
Proposition 2.4.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces with $F$ Dedekind complete and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a regular $m$-linear operator. If $A$ is disjointness preserving, then
$|A(|x_{1}|,\ldots,|x_{m}|)|=||A|(x_{1},\ldots,x_{m})|=|A(x_{1}\ldots,x_{m})|=|A|(|x_{1}|,\ldots,|x_{m}|)$
for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. Moreover,
$A^{+}(x_{1},\ldots,x_{m})=(A(x_{1},\ldots,x_{m}))^{+}\mbox{ and~{}}A^{-}(x_{1},\ldots,x_{m})=(A(x_{1},\ldots,x_{m}))^{-}$
for all $x_{1}\in E_{1}^{+},\ldots,x_{m}\in E_{m}^{+}$. In particular, $|A|$, $A^{+}$ and $A^{-}$ are Riesz multimorphisms.
Proof.
For simplicity, we prove the bilinear case $m=2$. Since $A$ is regular and $F$ is Dedekind complete, the modulus $|A|$ of $A$ exists and satisfies
$|A|(x_{1},x_{2})=\sup\left\{{\sum_{i_{1}=1}^{n_{1}}\sum_{i_{2}=1}^{n_{2}}|A(u_{i_{1}},v_{i_{2}})|}\,:\,{(u_{i_{1}})_{i_{1}=1}^{n_{1}}\in{\textstyle\prod_{x_{1}}},(v_{i_{2}})_{i_{2}=1}^{n_{2}}\in{\textstyle\prod_{x_{2}}}}\right\}$
for all $x_{1}\in E_{1}^{+},x_{2}\in E_{2}^{+}$ (see [21, Proposition 2.14]). Now, for all $a_{1}\in E_{1}^{+}$ and $a_{2}\in E_{2}^{+}$, the regular linear operators $A_{a_{1}}\colon E_{2}\longrightarrow F$ given by $A_{a_{1}}(x_{2})=A(a_{1},x_{2})$ and $A^{a_{2}}\colon E_{1}\longrightarrow F$ given by $A^{a_{2}}(x_{1})=A(x_{1},a_{2})$ preserve disjointness by assumption. By [5, Theorem 2.40] their moduli exist and
$|A_{a_{1}}|(|x_{2}|)=|A(a_{1},|x_{2}|)|=|A(a_{1},x_{2})|\text{ and }|A^{a_{2}}|(|x_{1}|)=|A(|x_{1}|,a_{2})|=|A(x_{1},a_{2})|$ (1)
for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Given $x_{1}\in E_{1}^{+}$, $x_{2}\in E_{2}^{+}$, $(u_{i})_{i=1}^{n_{1}}\in\prod_{x_{1}}$ and $(v_{j})_{j=1}^{n_{2}}\in\prod_{x_{2}}$, from (1) it follows that
$\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}|A(u_{i},v_{j})|$ $=\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}|A_{u_{i}}|(v_{j})=\sum_{i=1}^{n_{1}}|A_{u_{i}}|\left(\sum_{j=1}^{n_{2}}v_{j}\right)$
$=\sum_{i=1}^{n_{1}}|A_{u_{i}}|(x_{2})=\sum_{i=1}^{n_{1}}|A(u_{i},x_{2})|$
$=\sum_{i=1}^{n_{1}}|A^{x_{2}}|(u_{i})=|A^{x_{2}}|\left(\sum_{i=1}^{n_{1}}u_{i}\right)$
$=|A^{x_{2}}|(x_{1})=|A(x_{1},x_{2})|.$
This implies that $|A(x_{1},x_{2})|=|A|(x_{1},x_{2})$ for all $x_{1}\in E_{1}^{+}$ and $x_{2}\in E_{2}^{+}$. In particular, $|A|(|x_{1}|,|x_{2}|)=|A(|x_{1}|,|x_{2}|)|$ holds for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Therefore, using that $A(x_{1},x_{2}^{+})\perp A(x_{1},x_{2}^{-})$ and (1),
$|A(x_{1},x_{2})|$ $=|A(x_{1},x_{2}^{+})-A(x_{1},x_{2}^{-})|$
$=|A(x_{1},x_{2}^{+})+A(x_{1},x_{2}^{-})|=|A(x_{1},|x_{2}|)|=|A(|x_{1}|,|x_{2}|)|$
holds for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Then,
$|A|(x_{1},x_{2})$ $=|A|(x_{1}^{+},x_{2}^{+})-|A|(x_{1}^{+},x_{2}^{-})-|A|(x_{1}^{-},x_{2}^{+})+|A|(x_{1}^{-},x_{2}^{-})$
$=|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{+})|+|A(x_{1}^{-},x_{2}^{-})|$
for all for $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Since $A$ is disjointness preserving, $A(x_{1}^{+},x_{2}^{+})\perp A(x_{1}^{-},x_{2}^{+})$, so, by [5, Theorem 1.7(3)] we get
$\big{|}|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{-},x_{2}^{+})|\big{|}=|A(x_{1}^{+},x_{2}^{+})|\vee|A(x_{1}^{-},x_{2}^{+})|\leq|A|(|x_{1}|,x_{2}^{+}).$
Analogously, $\big{|}|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{-})|\big{|}\leq|A|(|x_{1}|,x_{2}^{-})$. Calling $a=|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{-},x_{2}^{+})|$ and $b=|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{-})|$, we obtain
$|a|\wedge|b|\leq|A|(|x_{1}|,x_{2}^{+})\wedge|A|(|x_{1}|,x_{2}^{-})=|A(x_{1},x_{2}^{+})|\wedge|A(x_{1},x_{2}^{-})|=0.$
By [5, Theorem 1.7],
$||A|(x_{1},x_{2})|$ $=|a-b|=|a+b|=|a|+|b|$
$=\big{|}|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{-},x_{2}^{+})|\big{|}+\big{|}|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{-})|\big{|}$
$=|A(x_{1}^{+},x_{2}^{+})|+|A(x_{1}^{-},x_{2}^{+})|+|A(x_{1}^{+},x_{2}^{-})|+|A(x_{1}^{-},x_{2}^{-})|$
$=|A|(x_{1}^{+},x_{2}^{+})+|A|(x_{1}^{-},x_{2}^{+})+|A|(x_{1}^{+},x_{2}^{-})+|A|(x_{1}^{-},x_{2}^{-})$
$=|A|(|x_{1}|,|x_{2}|),$
completing the proof of the desired formulas. The last statement follows from the formulas we have just proved. ∎
3 Arens extensions of disjointness preserving multilinear operators
We shall use the notation from [7] to introduce the Arens extensions of multilinear operators. Let $m\in\mathbb{N}$ and $\rho\in S_{m}$ be given. For $i=1,\ldots,m-1$, and $x_{\rho(i)}^{\prime\prime}\in E_{\rho(i)}^{\sim\sim}$, consider the operator
$\overline{x_{\rho(i)}^{\prime\prime}}\colon\mathcal{L}_{r}(E_{\rho(i)},\ldots,E_{\rho(m)})\longrightarrow\mathcal{L}_{r}(E_{\rho(i+1)},\ldots,E_{\rho(m)})~{},~{}\overline{x_{\rho(i)}^{\prime\prime}}(B)=x_{\rho(i)}^{\prime\prime}\circ B^{i},$
where $B^{i}\colon E_{\rho(i+1)}\times\cdots\times E_{\rho(m)}\longrightarrow E_{\rho(i)}^{\sim}$ is given by
$B^{i}(x_{i+1},\ldots,x_{m})(x_{i})=B(x_{i},x_{i+1},\ldots,x_{m})\text{ for all}x_{j}\in E_{\rho(j)},j=i,\ldots,m.$
In the case $i=m$, we define $\overline{x_{\rho(m)}^{\prime\prime}}\colon E_{\rho(m)}^{\sim}\longrightarrow\mathbb{R}$ by $\overline{x_{\rho(m)}^{\prime\prime}}=x_{\rho(m)}^{\prime\prime}$.
Given an $m$-linear form $C\colon E_{1}\times\cdots\times E_{m}\longrightarrow\mathbb{R}$, to each permutation $\rho\in S_{m}$ we consider the $m$-linear form
$C_{\rho}\colon E_{\rho(1)}\times\cdots\times E_{\rho(m)}\longrightarrow\mathbb{R}~{},~{}C_{\rho}(x_{1},\ldots,x_{m})=C(x_{\rho^{-1}(1)},\ldots,x_{\rho^{-1}(m)}).$
The Arens extension of a regular $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ with respect to the permutation $\rho$ is defined as the $m$-linear operator $AR_{m}^{\rho}(A)\colon E_{1}^{\sim\sim}\times\cdots\times E_{m}^{\sim\sim}\longrightarrow F^{\sim\sim}$ given by
$AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime})(y^{\prime})=\big{(}\overline{x_{\rho(m)}^{\prime\prime}}\circ\cdots\circ\overline{x_{\rho(1)}^{\prime\prime}}\big{)}((y^{\prime}\circ A)_{\rho})$
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\ldots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$ and $y^{\prime}\in F^{\sim}$. Although the notation is a bit different, these are exactly the Arens extensions studied in [6] and in many papers on Arens (or Aron-Berner) extensions. The extension $A^{(m+1)*}$ from [10, 12] is recovered by considering the permutation
$\theta(m)=1,\,\theta(m-1)=2,\ldots,\theta(2)=m-1,\,\theta(1)=m,$
that is, $AR_{m}^{\theta}(A)=A^{*(m+1)}$.
The first result on Arens extensions that preserve disjointness for bilinear operators is the following.
Proposition 3.1.
Let $E_{1},E_{2},F$ be Riesz spaces with $F$ Archimedean, and let $A\colon E_{1}\times E_{2}\longrightarrow F$ be a regular bilinear operator. If $A$ is disjointness preserving, then, for $\rho\in S_{2}$ and $x_{1}\in E_{1},x_{2}\in E_{2}$, the operators
$AR_{2}^{\rho}(A)(J_{E_{1}}(x_{1}),\cdot)\colon E_{2}^{\sim\sim}\longrightarrow F^{\sim\sim},\,AR_{2}^{\rho}(A)(J_{E_{1}}(x_{1}),\cdot)(y_{2}^{\prime\prime})=AR_{2}^{\rho}(A)(J_{E_{1}}(x_{1}),y_{2}^{\prime\prime})$
and
$AR_{2}^{\rho}(A)(\cdot,J_{E_{2}}(x_{2}))\colon E_{1}^{\sim\sim}\longrightarrow F^{\sim\sim},\,AR_{2}^{\rho}(A)(\cdot,J_{E_{2}}(x_{2}))(x_{2}^{\prime\prime})=AR_{2}^{\rho}(A)(x_{1}^{\prime\prime},J_{E_{2}}(x_{2})),$
are disjointness preserving.
Proof.
In this case we only have two permutations, namely $\theta$ ($\theta(1)=2,\theta(2)=1$) and $id$ ($id(1)=1,id(2)=2)$. Let us check that $AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1}),\cdot)=A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving for every $x_{1}\in E_{1}$. Indeed, for all $x_{1}\in E_{1},\,x_{2}^{\prime\prime}\in E_{2}^{\sim\sim}$ and $y^{\prime}\in F^{\sim}$,
$A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),x_{2}^{\prime\prime})(y^{\prime})$ $=J_{E_{1}}(x_{1})\big{(}\overline{x_{2}^{\prime\prime}}^{\theta}(y^{\prime}\circ A)\big{)}=\overline{x_{2}^{\prime\prime}}^{\theta}(y^{\prime}\circ A)(x_{1})$
$=x_{2}^{\prime\prime}(y^{\prime}\circ A(x_{1},\cdot))=x_{2}^{\prime\prime}(A(x_{1},\cdot)^{\prime}(y^{\prime}))=A(x_{1},\cdot)^{\prime\prime}(x_{2}^{\prime\prime})(y^{\prime}),$
hence $A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),x_{2}^{\prime\prime})=A(x_{1},\cdot)^{\prime\prime}(x_{2}^{\prime\prime})$ for all $x_{1}\in E_{1}$ and $x_{2}^{\prime\prime}\in E_{2}^{\sim\sim}$. Since $A(x_{1},\cdot)$ is disjointness preserving by assumption, so is its second adjoint $A(x_{1},\cdot)^{\prime\prime}$ by Theorem 2.1(1). Therefore, $A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving. The same argument shows that $AR_{2}^{id}(A)(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving.
Now, we check that $AR_{2}^{\theta}(A)(\cdot,J_{E_{2}}(x_{2}))=A^{\ast\ast\ast}(\cdot,J_{E_{2}}(x_{2}))$ is also disjointness preserving. First, observe that for all $x_{1}\in E_{1},x_{2}\in E_{2}$ and $y^{\prime}\in F^{\sim}$,
$\overline{J_{E_{2}}(x_{2})}^{\theta}(y^{\prime}\circ A)(x_{1})=J_{E_{2}}(x_{2})(y^{\prime}\circ A(x_{1},\cdot))=y^{\prime}(A(x_{1},x_{2}))=[y^{\prime}\circ A(\cdot,x_{2})](x_{1}),$
which implies that $\overline{J_{E_{2}}(x_{2})}^{\theta}(y^{\prime}\circ A)=y^{\prime}\circ A(\cdot,x_{2})$. Thus, given $x_{1}^{\ast\ast}\in E_{1}^{\ast\ast}$, $x_{2}\in E_{2}$ and $y^{\prime}\in F^{\sim}$, we have
$A^{\ast\ast\ast}(x_{1}^{\prime\prime},J_{E_{2}}(x_{2}))(y^{\prime})$ $=x_{1}^{\prime\prime}\big{(}\overline{J_{E_{2}}(x_{2})}^{\theta}(y^{\prime}\circ A)\big{)}=x_{1}^{\prime\prime}(y^{\prime}\circ A(\cdot,x_{2}))$
$=x_{1}^{\prime\prime}(A(\cdot,x_{2}))^{\prime}(y^{\prime}))=A(\cdot,x_{2})^{\prime\prime}(x_{1}^{\prime\prime})(y^{\prime}).$
This proves that $A^{\ast\ast\ast}(x_{1}^{\prime\prime},J_{E_{2}}(x_{2}))=A(\cdot,x_{2})^{\prime\prime}(x_{1}^{\prime\prime})$ for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim}$ and $x_{2}\in E_{2}$. Since $A(\cdot,x_{2})$ is disjointness preserving by assumption, we get from Theorem 2.1(1) that its second adjoint is also disjointness preserving, which implies that $A^{\ast\ast\ast}(\cdot,J_{E_{2}}(x_{2}))$ is disjointness preserving as well. The same argument shows that $AR_{2}^{id}(A)(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving. ∎
As in [6, p. 11], a map taking values in a Riesz space is said to have finite rank if the sublattice generated by its range is finite dimensional. Now, we prove our first main result, which shows that all Arens extensions of a finite rank regular disjointness preserving multilinear operator into a Dedekind complete Riesz space are also disjointness preserving.
Theorem 3.2.
Let $E_{1},\ldots,E_{m},F$ be Riesz space with $F$ Dedekind complete and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a finite rank regular $m$-linear operator. If $A$ is disjointness preserving, then all Arens extensions $AR_{m}^{\rho}(A)$ of $A$, $\rho\in S_{m}$, are disjointness preserving.
Proof.
We split the proof into three steps. First we prove the case $F=\mathbb{R}$. Given a regular disjointness preserving $m$-linear form $A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$, from Proposition 2.4 we know that $|A|$ is a scalar-valued Riesz multimorphism. By [16, Theorem 6] there are scalar-valued Riesz homomorphisms $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{m}\in E_{m}^{\sim}$ such that $|A|(x_{1},\ldots,x_{m})=\varphi_{1}(x_{1})\cdots\varphi_{m}(x_{m})$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. By the proof of [6, Theorem 3.2],
$AR_{m}^{\rho}(|A|)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime})=\varphi_{1}^{\prime\prime}(x_{1}^{\prime\prime})\cdots\varphi_{m}^{\prime\prime}(x_{m}^{\prime\prime})$
holds for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\ldots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$ and $\rho\in S_{m}$. Since $\varphi_{1},\dots,\varphi_{m}$ are disjointness preserving, their second adjoints $\varphi_{1}^{\prime\prime},\dots,\varphi_{1}^{\prime\prime}$ are disjointness preserving by Theorem 2.1(1). Then, $AR_{m}^{\rho}(|A|)$ is disjointness preserving. By [6, Theorem 2.2(c)] we know that the correspondence
$A\in\mathcal{L}_{r}(E_{1},\ldots,E_{m};F)\mapsto AR_{m}^{\rho}(A)\in\mathcal{L}_{r}(E_{1}^{\sim\sim},\ldots,E_{m}^{\sim\sim};F^{\sim\sim})$
is a positive linear operator. So, $A\leq|A|$ implies $AR_{m}^{\rho}(A)\leq AR_{m}^{\rho}(|A|)$, which, together with the disjointness preservation of $AR_{m}^{\rho}(|A|)$, gives that $AR_{m}^{\rho}(A)$ is also disjointness preserving.
In the case $F=\mathbb{R}^{n},n\in\mathbb{N}$, we apply the previous case to the compositions $\pi_{j}\circ A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}^{n}$, where $\pi_{j}\colon\mathbb{R}^{n}\to\mathbb{R}$ is the projection in the $j$-th coordinate, $j=1,\ldots,n$, to obtain that $AR_{m}^{\rho}(\pi_{j}\circ A)\colon E_{1}^{\sim\sim}\times\cdots\times E_{m}^{\sim\sim}\longrightarrow\mathbb{R}$ is disjointness preserving for all $\rho\in S_{m}$ and $j=1,\dots,n$. By [6, Remark 2.3],
$\pi_{j}\circ AR_{m}^{\rho}(A)=\pi_{j}^{\prime\prime}\circ AR_{m}^{\rho}(A)=AR_{m}^{\rho}(\pi_{j}\circ A)$
holds for all $\rho\in S_{m}$ and $j=1,\dots,m$. Therefore,
$AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime})$ $=(\pi_{1}\circ AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime}),\dots,\pi_{n}\circ AR_{m}^{\rho}(A))(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime}))$
$=(AR_{m}^{\rho}(\pi_{1}\circ A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime}),\dots,AR_{m}^{\rho}(\pi_{n}\circ A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime}))$
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$. As $\mathbb{R}^{n}$ is endowed with the coordinatewise order, it follows that $AR_{m}^{\rho}(A)$ is disjointness preserving for every $\rho\in S_{m}$.
Finally, let $F$ be a Dedekind complete Riesz space and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a finite rank disjointness preserving regular operator. Calling $G$ be the finite-dimensional sublattice of $F$ generated by the range of $A$, by $i\colon G\longrightarrow F$ we denote the inclusion operator. It is clear that $A_{1}\colon E_{1}\times\cdots\times E_{m}\longrightarrow G,A_{1}(x_{1},\ldots,x_{m})=A(x_{1},\ldots,x_{m})$, is a disjointness preserving $m$-linear operator. Since $G$ is Archimedean, by [24, Corollary 1, p. 70] there exists a Riesz isomorphism $I\colon G\to\mathbb{R}^{n}$, where $n=\dim G$ and $\mathbb{R}^{n}$ is endowed with the coordinatewise order. Thus, $I\circ A_{1}$ is an $\mathbb{R}^{n}$-valued disjointness preserving multilinear operator, so $AR_{m}^{\rho}(I\circ A_{1})$ is disjointness preserving by the second part of this proof. From [6, Remark 2.3] we have
$AR_{m}^{\rho}(A)=AR_{m}^{\rho}(i\circ I^{-1}\circ I\circ A_{1})=i^{\prime\prime}\circ(I^{-1})^{\prime\prime}\circ AR_{m}^{\rho}(I\circ A_{1}).$
Therefore, $AR_{m}^{\rho}(A)$ is disjointness preserving because $i^{\prime\prime}$ and $(I^{-1})^{\prime\prime}$ are disjointness preserving linear operators by Theorem 2.1. ∎
Now we proceed to the case of arbitrary regular multilinear operators. The following two results shall be helpful.
Proposition 3.3.
Let $E_{1},\dots,E_{m},\,F$ be Riesz spaces, let $A\colon E_{1}\times\cdots\times E_{m}\to F$ be a regular disjointness preserving $m$-linear operator, and let $\rho\in S_{m}$ be given. If $y^{\prime}\in F^{\sim}$ is disjointness preserving, then $J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ is disjointness preserving and
$|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|(|y^{\prime}|)=|AR_{m}^{\rho}(A)(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)(y^{\prime})|=|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})(y^{\prime})|$
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$.
Proof.
Since $A$ is regular and disjointness preserving and $y^{\prime}\in F^{\sim}$ is disjointness preserving, $y^{\prime}\circ A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$ is a regular disjointness preserving $m$-linear operator. By the proof of Theorem 3.2 we get that $AR_{m}^{\rho}(y^{\prime}\circ A)$ is also disjointness preserving. The proof of [6, Proposition 3.4] yields that $AR_{m}^{\rho}(y^{\prime}\circ A)=J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ holds, proving that $J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ is disjointness preserving and that
$|J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)|=|J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|$ (2)
holds for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$. Observe also that $J_{F^{\sim}}(y^{\prime})$ is disjointness preserving by Theorem 2.1(1), hence
$|J_{F^{\sim}}(y^{\prime})|(|y^{\prime\prime}|)=|J_{F^{\sim}}(y^{\prime})(y^{\prime\prime})|$ (3)
for every $y^{\prime\prime}\in F^{\sim\sim}$. Combining (2) and (3), we have
$|AR_{m}^{\rho}(A)(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)(y^{\prime})|$ $=|(J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A))(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)|$
$=|(J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A))(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|$
$=|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})(y^{\prime})|$
$\leq|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|(|y^{\prime}|)$
$=J_{F^{\sim}}(|y^{\prime}|)(|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|)$
$=|J_{F^{\sim}}(y^{\prime})|(|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime})|)$
$=|J_{F^{\sim}}(y^{\prime})(AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\dots,x_{m}^{\prime\prime}))|$
$=|AR_{m}^{\rho}(A)(|x_{1}^{\prime\prime}|,\dots,|x_{m}^{\prime\prime}|)(y^{\prime})|$
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$. ∎
Next result can be regarded as a version of [6, Proposition 3.6] for disjointness preserving multilinear operators.
Proposition 3.4.
Let $E_{1},\ldots,E_{m},F$ be Banach lattices, let $A\in{\cal L}_{r}(E_{1},\ldots,E_{m};F)$ be a disjointness preserving operator, and let $\rho\in S_{m}$ be given. Then:
(a) $y^{***}\circ AB_{m}^{\rho}(A)$ is a disjointness preserving $m$-linear operator for every weak*-continuous disjointness preserving functional $y^{***}\in F^{***}$.
(b) Fix $i\in\{1,\ldots,m\}$ and let $w_{i}^{\ast\ast},z_{i}^{\ast\ast}\in E_{i}^{\ast\ast}$ be such that $w_{i}^{\ast\ast}\perp z_{i}^{\ast\ast}$. Then
$AB_{m}^{\rho}(A)(x_{1}^{**},\ldots,w_{i}^{\ast\ast},\ldots,x_{m}^{**})(y^{*})\perp AB_{m}^{\rho}(A)(x_{1}^{**},\ldots,z_{i}^{\ast\ast},\ldots,x_{m}^{**})|)(y^{*})\mbox{ in }\mathbb{R}$
for all $x_{j}^{**}\in E_{j}^{**},j\neq i$ and any $y^{*}\in\overline{{\rm span}}\{\varphi\in F^{\ast}:\varphi\text{ is disjointness preserving}\}.$
Proof.
(a) Let $y^{***}\in F^{***}$ be a weak*-continuous disjointness preserving functional, and let $y^{\ast}\in F^{\ast}$ be such that $y^{\ast\ast\ast}=J_{F^{\ast}}(y^{\ast})$. As $y^{\ast\ast\ast}$ is disjointness preserving and $J_{F}$ is a Riesz homomorphisms, $y^{\ast}=y^{\ast\ast\ast}\circ J_{F}$ is disjointness preserving. Now, Proposition 3.3 yields that $y^{***}\circ AB_{m}^{\rho}(A)=J_{F^{\ast}}(y^{\ast})\circ AB_{m}^{\rho}(A)$ is a disjointness preserving operator.
(b) For each $j\neq i$, take $x_{j}^{\ast\ast}\in E_{j}^{\ast\ast}$. Given $y^{\ast}\in\overline{{\rm span}}\{\varphi\in F^{\ast}:\varphi\text{ is a disjointness preserving}\}$, there exists a sequence $(y_{n}^{\ast})_{n=1}^{\infty}$ in ${\rm span}\{\varphi\in F^{\ast}:\varphi\text{ is a disjointness preserving}\}$ such that $y_{n}^{\ast}\longrightarrow y^{\ast}$ in $F^{*}$, hence $|y_{n}^{\ast}|\longrightarrow|y^{\ast}|$. Moreover, for each $n\in\mathbb{N}$ there exist disjointness preserving linear functionals $\varphi^{1}_{n},\ldots,\varphi_{n}^{k_{n}}\in F^{\ast}$ and scalars $\alpha_{n}^{1},\ldots,\alpha_{n}^{k_{n}}$ such that $y_{n}^{\ast}=\sum\limits_{j=1}^{k_{n}}\alpha_{n}^{j}\varphi_{n}^{j}$. For each $w\in E_{i}^{\ast\ast}$, write
$AB_{m}^{\rho}(A)_{i}(w):=AB_{m}^{\rho}(A)(x_{1}^{\ast\ast},\ldots,x_{i-1}^{\ast\ast},w,x_{i+1}^{\ast\ast},\ldots,x_{m}^{\ast\ast}).$
It follows from Proposition 3.3 that $|AB_{m}^{\rho}(A)_{i}(w)|(|\varphi|)=|AB_{m}^{\rho}(A)_{i}(w)(\varphi)|$ for every disjointness preserving $\varphi\in F^{\ast}$ and every $w\in E_{i}^{\ast\ast}$. Therefore,
$\big{(}|AB_{m}^{\rho}$ $(A)_{i}(w_{i}^{\ast\ast})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|\big{)}(|y^{*}|)$
$=\lim_{n\rightarrow\infty}\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|\big{)}(|y_{n}^{\ast}|)$
$=\lim_{n\rightarrow\infty}\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|\big{)}\Big{(}\Big{|}\sum\limits_{j=1}^{k_{n}}\alpha_{n}^{j}\varphi_{n}^{j}\Big{|}\Big{)}$
$\leq\lim_{n\rightarrow\infty}\displaystyle\sum_{j=1}^{k_{n}}|\alpha_{n}^{j}|\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|\big{)}(|\varphi_{n}^{j}|)$
$\leq\lim_{n\rightarrow\infty}\displaystyle\sum_{j=1}^{k_{n}}|\alpha_{n}^{j}|\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})|(|\varphi_{n}^{j}|)\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})|(|\varphi_{n}^{j}|)\big{)}$
$=\lim_{n\rightarrow\infty}\displaystyle\sum_{j=1}^{k_{n}}|\alpha_{n}^{j}|\big{(}|AB_{m}^{\rho}(A)_{i}(w_{i}^{\ast\ast})(\varphi_{n}^{j})|\wedge|AB_{m}^{\rho}(A)_{i}(z_{i}^{\ast\ast})(\varphi_{n}^{j})|\big{)}$
$=\lim_{n\rightarrow\infty}\displaystyle\sum_{j=1}^{k_{n}}|\alpha_{n}^{j}|\big{(}|(J_{F^{\ast}}(\varphi_{n}^{j})\circ AB_{m}^{\rho}(A)_{i})(w_{i}^{\ast\ast})|\wedge|(J_{F^{\ast}}(\varphi_{n}^{j})\circ AB_{m}^{\rho}(A)_{i})(z_{i}^{\ast\ast})|\big{)}=0,$
where the last equality follows from Proposition 3.3. ∎
Our second main result is a consequence of Proposition 3.4(b).
Corollary 3.5.
Let $F$ be a Dedekind complete Banach lattice such that $F^{\ast}$ has a Schauder basis consisting of disjointness preserving linear functionals. Then all Arens extensions of any $F$-valued disjointness preserving regular multilinear operator are disjointness preserving.
Example 3.6.
Corollary 3.5 holds for the following choices of $F$: $c_{0}$, $\ell_{p}$ with $1<p<\infty$, Tsirelson’s original space $T^{\ast}$ and its dual $T$, Schreier’s space $S$, and the predual $d_{\ast}(w,1)$ of the Lorentz sequence space $d(w,1)$ (see [6, Example 3.9]).
From here onwards, we shall use that every Banach lattice is Archimedean [22, Proposition 1.1.8(i)]. Recall that $\theta$ is the permutation of $\{1,2\}$ given by $\theta(1)=2,\theta(2)=1$.
Proposition 3.7.
Let $E_{1},E_{2},F$ be Banach lattices such that $E_{1}^{*}$ is separable and the lattice operations in $F^{**}$ are weak*-sequentially continuous. Let $A\colon E_{1}\times E_{2}\longrightarrow F$ be a regular bilinear operator. If $A$ is disjointness preserving, then the operator
$AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},\cdot)\colon E_{2}^{\ast\ast}\longrightarrow F^{\ast\ast},\,AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},\cdot)(x_{2}^{\ast\ast})=AR_{2}^{\rho}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast}),$
is disjointness preserving for each $x_{1}^{\ast\ast}\in E_{1}^{\ast\ast}$.
Proof.
Let $x_{2}^{\ast\ast},z_{2}^{\ast\ast}\in E_{2}^{\ast\ast}$ be such that $x_{2}^{\ast\ast}\perp z_{2}^{\ast\ast}$. By Proposition 3.1, we have
$AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1}),x_{2}^{\ast\ast})\perp AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1}),z_{2}^{\ast\ast})\text{ for all }x_{1}\in E_{1}.$
Let $x_{1}^{**}\in E_{1}^{**}$ be such that $\|x_{1}^{**}\|=1$. Since $E_{1}^{*}$ is separable, the weak topology of $E_{1}^{**}$ is metrizable on the bounded sets of $E_{1}^{**}$. So, as $J(B_{E_{1}})$ is weak-dense in $B_{E_{1}^{**}}$, there exists a sequence $(x_{1,n})_{n}$ in $B_{E_{1}}$ such that $J(x_{1,n})\xrightarrow{\,\omega^{\ast}}x_{1}^{**}$ in $E_{1}^{**}$. Since the Arens extension $AR_{2}^{\theta}$ is weak continuous in the first variable [13, p. 413], $AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},z_{2}^{\ast\ast})=\omega^{\ast}-\displaystyle\lim_{n}AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),z_{2}^{\ast\ast})$ and $AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast})=\omega^{\ast}-\displaystyle\lim_{n}AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),x_{2}^{\ast\ast})$. Thus,
$|AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast})|$ $\wedge|AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},z_{2}^{\ast\ast})|$
$=\omega^{\ast}-\lim_{n}|AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),x_{2}^{\ast\ast})|\wedge\omega^{\ast}-\lim_{n}|AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),z_{2}^{\ast\ast})|$
$=\omega^{\ast}-\lim_{n}(|AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),x_{2}^{\ast\ast})|\wedge|AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),z_{2}^{\ast\ast})|)=0.$
Thus far we have $AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast})\perp AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},z_{2}^{\ast\ast})$ for all $x_{1}^{**}\in E_{1}^{**}$ with $\|x_{1}^{**}\|=1$. For $0\neq z_{1}^{**}\in E_{1}^{**}$, $w_{1}^{\ast\ast}:=\frac{z_{1}^{**}}{\|z_{1}^{**}\|}$ satisfies $\|w_{1}^{**}\|=1$ and $AR_{2}^{\theta}(A)(w_{1}^{\ast\ast},x_{2}^{\ast\ast})\perp AR_{2}^{\theta}(A)(w_{1}^{\ast\ast},z_{2}^{\ast\ast})$. It follows that $AR_{2}^{\theta}(A)(z_{1}^{\ast\ast},x_{2}^{\ast\ast})\perp AR_{2}^{\theta}(A)(z_{1}^{\ast\ast},z_{2}^{\ast\ast})$, which proves that the operator $AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},\cdot)$ is disjointness preserving. ∎
Using that the Arens extension $AR_{2}^{id}(A)\colon E_{1}^{\ast\ast}\times E_{2}^{\ast\ast}\longrightarrow F^{\ast\ast}$ is weak continuous in the second variable and Proposition 3.1, the same proof of Proposition 3.7 gives the following.
Proposition 3.8.
Let $E_{1},E_{2},F$ be Banach lattices such that $E_{2}^{*}$ is separable and the lattice operations in $F^{**}$ are weak*-sequentially continuous. Let $A\colon E_{1}\times E_{2}\longrightarrow F$ be a regular bilinear operator. If $A$ is disjointness preserving, then the operator
$AR_{2}^{id}(A)(\cdot,x_{2}^{\ast\ast})\colon E_{1}^{\ast\ast}\longrightarrow F^{\ast\ast},\,AR_{2}^{id}(A)(\cdot,x_{2}^{\ast\ast})(x_{1}^{\ast\ast})=AR_{2}^{id}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast}),$
is disjointness preserving for each $x_{2}^{\ast\ast}\in E_{2}^{\ast\ast}$.
Just combine Propositions 3.7 and 3.8 to get the following.
Corollary 3.9.
Let $E_{1},E_{2}$ be Banach lattices with separable duals, and let $F$ be a Banach lattice such that the lattice operations in $F^{**}$ are weak*-sequentially continuous. Let $A\colon E_{1}\times E_{2}\longrightarrow F$ be a regular disjointness preserving bilinear operator whose Arens extensions coincide. Then the (unique) Arens extensions of $A$ is disjointness preserving.
Recall that a Banach space $E$ is Arens regular if every bounded linear operator from $E$ to $E^{\ast}$ is weakly compact. By [26, Theorem 2.2], the Banach space $E$ is Arens regular if and only if the Arens extensions of every continuous bilinear operator from $E^{2}$ to any Banach space are coincident.
Corollary 3.10.
Let $E$ be an Arens regular Banach lattice with separable dual, and let $F$ be a Banach lattice such that the lattice operations in $F^{**}$ are weak*-sequentially continuous. Then, the unique Arens extension of any disjointness preserving regular $m$-linear operator from $E^{2}$ to $F$ is disjointness preserving.
Examples of nonreflexive Arens regular Banach lattices are $c_{0},\ell_{\infty}$ and $C(K)$ for any infinite compact Hausdorff space $K$. In particular, AM-spaces with order unit are Arens regular.
Remark 3.11.
The last four results can be proved, mutatis mutandis, for disjointness preserving regular $m$-linear operators, $m\geq 2$. The reason is that every Arens extension of a continuous $m$-linear operator is weak continuous in one variable. For details, see the proof of [6, Lemma 2.4].